The width of a rectangular parking lot is 8 feet less than three-fifths of the length. The perimeter of the lot is 400 feet. Find the length and width of the lot.
Answer by mananth(16946)

length =x
width = (3/5)x-8

Perimeter = 2(L+W)
=2(x+(3/5)x-8))
400=2(x+(3/5)x-8))
/2
200 = x+(3/5)x-8
200-8= x+(3/5)x
192=(8/5)x
192*5=8x
192*5/8 =x
24*5=x
x=120
length =120